OpenStudy (anonymous):

Solve the equation by completing the square. x^2+3x-9=0 Will fan & Medal

OpenStudy (janu16):

x2+3x−9=0 : x=3(√5−1)/2 , x=−3(1+√5)/2

OpenStudy (janu16):

This / means over 2

OpenStudy (janu16):

Do you want the steps? @arteryx

OpenStudy (anonymous):

@Janu16 Yes, please.